THE MICROWAVE SPECTRUM OF PARTIALLY DEUTERATED SPECIES OF DIMETHYL ETHER^{a}

D. LAUVERGNAT, Laboratoire de Chimie Physique, Bât. 349, CNRS, UMR8000, Université Paris-Sud, Orsay, F-91405, France; L. MARGULÈS, R. A. MOTIYENKO, Laboratoire PhLAM, CNRS/Université des Sciences et Technologies de Lille 1, Bât. P5, 59655 Villeneuve d’Ascq, France; J.-C. GUILLEMIN, Sciences Chimiques de Rennes, UMR6226 CNRS-ENSCR, Avenue du Général Leclerc, CS 50837, 35708 Rennes Cedex 7, France; AND \underline{L. H. COUDERT}, LISA, CNRS/Universités Paris Est et Paris Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil, France.

Dimethyl ether is a molecule of astrophysical interest^{b} spectroscopically well characterized.^{c} It is one of the simplest molecules with two methyl groups undergoing large amplitude internal rotations. Due to deuterium enrichment in the interstellar medium,^{d} one can reasonably expect that partially deuterated species of dimethyl ether might be detected. However, there are no spectroscopic results about the microwave spectrum of such species. A theoretical calculation of the rotation-torsion energy levels of the partially deuterated species of dimethyl ether has been undertaken aided by ab initio calculations. The approach accounts for the complicated torsion-rotation interactions displayed by this molecule and for the fact that deuteration leads to changes of the bidimensional internal rotation effective potential energy surface. Due to zero-point energy contributions from the 19 small amplitude vibrational modes, this surface no longer displays G_{36} symmetry. Rotation-torsion energy levels are computed treating the two angles of internal rotation as active coordinates and evaluating Hamiltonian matrix elements with the help of Gaussian quadrature.^{e} It is hoped that the present results will allow us to understand the microwave spectrum of the mono deuterated species CH_{2}DOCH_{3} which has been recorded in Lille with the new sub millimeter wave spectrometer (150–950 GHz) based on harmonic generation of solid-state sources.

aThis work is supported by ANR-08-BLAN-0054, ANR-08-BLAN-0225, and by the PCMI French program.

bSnyder, Buhl, and Schwartz, Astrophys. J. Letters 191 (1974) L79.

cEndres, Drouin, Pearson, Müller, Lewen, Schlemmer, and Giesen, A&A 504 (2009) 635.

dSolomon and Woolf, Astrophys. J. Letters 180 (1973) L89.

eLauvergnat and Nauts, J. Chem. Phys. 116 (2002) 8560; and Light and Bačić, J. Chem. Phys. 87 (1987) 4008.